OpenStudy (anonymous):

Please help For every 3 laps that Jackie swims, Ronnie swims 2 laps. During practice, they swam a combined total of 70 laps. How many laps did Ronnie swim during practice? 42 36 28 24

OpenStudy (anonymous):

I know I will tell u the answer

OpenStudy (anonymous):

Okay? Can you walk me through how to do it, so I know how to do it on other problems?

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

Okay. I appreciate it.

OpenStudy (anonymous):

I am in 6th grade but I will try to figure it out

OpenStudy (agl202):

28 is correct answer, people!!!

OpenStudy (anonymous):

Oh, I'm a freshmen. You will probably have a hard time......... Sorry.....

OpenStudy (anonymous):

they say that jakie swam 3 laps, and we know that both and ronnie swam a total of 70 laps so I think it is 28

OpenStudy (anonymous):

Thank you @Agl202 but how did you get that answer?

OpenStudy (campbell_st):

ok... so this is a ratio question... so use the unitary method add the parts in the ratio 3 + 2 = 5 so jackie = 3/5 x 70 ronnie = 2/5 x 70

OpenStudy (agl202):

3X12=42 FOR Jackie, 14X2 for Ronnie=70 Total!!! Not that hard at all!=)

OpenStudy (anonymous):

Uh. Okay. Thank you! I'm really confused, but I appreciate your assistance :)

OpenStudy (anonymous):

3 + 2 = 5 ronnie = 2/5 x 70 so jackie = 3/5 x 70

OpenStudy (anonymous):

is that helpful

OpenStudy (anonymous):

Yes, very!

OpenStudy (anonymous):

okay can u give me a medal

OpenStudy (anonymous):

I already did.

OpenStudy (anonymous):

okay hope this is helpful to u

OpenStudy (anonymous):

Lol, I always give a medal to whoever helps me understand.

OpenStudy (anonymous):

I will give u a medal to for working hard

OpenStudy (anonymous):

Thank you!! :)